OpenStudy (anonymous):

Factor completely: 18x2 - 21x - 15 3(6x + 1)(x - 5) 3(2x - 1)(3x + 5) 3(2x + 1)(3x - 5) 3(2x - 5)(3x + 1)

OpenStudy (anonymous):

First factor 3 out \[ 18 x^2 - 21 x - 15= 3(6x^2 -7x -5) \]

OpenStudy (anonymous):

\[ 18 x^2-21 x-15=3 (2 x+ ?) (3 x-?) \]

OpenStudy (anonymous):

\[ 18 x^2-21 x-15=3 (2 x+1) (3 x-5) \]

OpenStudy (anonymous):

Thanks for explaining it, I was just about to type that answer!

OpenStudy (anonymous):

That means you understood what was involved. Great.